OpenStudy (anonymous):

how do you simplify 2x^3 + 5x^2-3x/ 2x^2- 7x+ 3

OpenStudy (anonymous):

2x^2 - 3x^2 = -x^2 5x - -7x = 5x + 7x = 12x -4 - 1 = -5 -x^2 + 12x - 5

OpenStudy (anonymous):

first u must factories both the top part and the bottom part top 2x^3 + 5x^2 - 3x = x(2x-1)(x+3) bottom 2x^2 - 7x +3 = (2x-1)(x-3)

OpenStudy (anonymous):

then u cut off 2x-1 and its simplified

OpenStudy (anonymous):

ooh ok now i get this thank you!

OpenStudy (anonymous):

=]